The angle of diffracted X-rays is 90 degree from the incident direction of X-rays. The interplanner seperation of the planes responsible for the diffraction is?
1
Expert's answer
2020-06-17T09:35:00-0400

According to Bragg's law for n=1:

"\lambda=2d\text{ sin}\theta."

Hence, the interplanar spacing is


"d=\frac{\lambda}{2\text{sin}\theta}=\frac{\lambda}{2}."